# XRF V2: A Dataset for Action Summarization with Wi-Fi Signals, and IMUs in Phones, Watches, Earbuds, and Glasses

### v1

@@ 3.1. Hardware Setups
Wi-Fi transceivers. The Wi-Fi transceiver devices include one transmitter and three receivers, positioned at the four corners of a 3.1 m $\times$ 2.4 m rectangular area. This rectangle defines the activity region for human actions, and the devices are placed at a height of 120 cm to ensure effective coverage for full-body motion capture. The transmitter operates with a single antenna, broadcasting packets at 200 per second on channel 128 (5.64 GHz). Each receiver is equipped with three antennas. Using the Wi-Fi CSI Tool (Halperin et al., 2011), we extract the channel state information (CSI) from 30 orthogonal frequency-division multiplexing (OFDM) subcarriers. The resulting CSI recordings are represented as a tensor of size $(200t)\times 1\times 3\times 3\times 30$, where $t$ is the recording duration in seconds. For example, for an 80-second continuous action, the CSI recordings yield a tensor of size $16000\times 1\times 3\times 3\times 30$, capturing detailed variations in the wireless channel over time for subsequent analysis.
@@
Azure Kinect. The Azure Kinect captures RGB, depth, and infrared images at 30 frames per second in 720P resolution using its toolkit. The Kinect is placed between the transmitter and the third receiver at a height of 115 cm to capture video footage of the movement of individuals in the scene. We can use open-source computer vision models to extract pose and mesh information from the videos, which will be useful for future tasks such as IMU- and Wi-Fi-based pose estimation and mesh reconstruction.
The experimental scenarios are shown in Fig. 2. Before data recording, all the aforementioned hardware devices undergo careful synchronization. Notably, XRF V2 selects commonly available consumer-grade products used in everyday household and personal settings, unlike XRF55 (Wang et al., 2024a), which utilizes millimeter-wave radar and an RFID reader with a tag array.
@@
3.2. Action Sequence Proposal
We select three typical household scenarios for the experiment: the dining room, study room, and bedroom. These settings allow us to summarize users’ activities related to eating, studying, and sleeping, along with associated series of actions. For example, activities such as taking medicine, drinking water, washing hands, and stretching serve as health-related indicators; and actions like opening curtains or turning on lights provide insights into interactions with household devices.
We communicate the above objective to the volunteers, but instead of assigning each person a fixed action sequence, we aim to increase the diversity of action sequences and incorporate personal habits relevant to the volunteers. Therefore, we ask each volunteer to recall their own daily action sequences in the dining room, study room, and bedroom, and estimate the execution time of each action in the sequence. Each volunteer recalls five action sequences based on the starting positions of the actions, with each sequence containing 7-10 actions. Fig. 4 shows examples of some action sequences, while Table 2 presents the proposed actions in each scene.
@@ 3.3. Volunteer Coordination and Action Execution
In general, the duration of a single action sequence is approximately 50–80 seconds. After completing one action sequence, volunteers rest for 10 seconds, and after executing five sequences, they take a 5-minute break.
@@ 3.4. Dataset Statistics
Each of the 16 volunteers proposed 5 action sequences for each scene. By shortening durations and adjusting the length of individual actions, we generated multiple variants of these sequences. Specifically, in the bedroom, each volunteer performed three different variations of each action sequence, while in the study room and dining room, four different variations were performed for each sequence. Thus, XRF V2 dataset contains a total of: $16\times 5\times 3+16\times 5\times 4+16\times 5\times 4=880$ action sequences. After filtering out incomplete and erroneous sequences, we obtained 853 valid action sequences, with a total duration of 16 hours, 16 minutes and 8 seconds. These sequences were captured using synchronized Channel State Information (CSI) from four Wi-Fi transceivers, IMU data from two smartwatches, two smartphones, a pair of earbuds, and a pair of smart glasses, along with RGB+D+IR video streams from a Kinect camera. Furthermore, we implemented a stratified partitioning strategy maintaining an 80:20 ratio between training and testing sets across all environmental contexts and participants, resulting in 682 sequences (80%) allocated for model training and 171 sequences (20%) reserved for evaluation purposes. Table. 3 shows detailed statistics of XRF V2 dataset.
@@
4. Method
State Space Models (SSMs), such as S4 (Gu et al., 2021), S5 (Smith et al., 2022), and Mamba (Gu and Dao, 2023), are a recent class of deep learning methods that have demonstrated advantages in both natural language processing (Gu and Dao, 2023) and computer vision (Zhu et al., 2024). Unlike Transformer-based models, which suffer from quadratic time complexity due to self-attention, SSMs operate with linear-time complexity, making them significantly more efficient for long sequence modeling. In action summarization, the input sensor sequences are much longer than those used in action recognition tasks, necessitating the ability to effectively model long-range dependencies. To this end, we base our approach on an advanced SSM model and propose the XRFMamba network. Next, we first introduce the preliminary concepts of SSMs in Sec. 4.1, then, we will present XRFMamba in Sec. 4.2.
4.1. State Space Model Preliminaries
@@
Due to the flexibility of SSMs, they can switch to convolution during training and to recurrence during inference, providing fast performance in both stages. SSMs have a rich foundational knowledge, and introducing all aspects is beyond the scope of this paper. For more detailed information, please refer to Mamba (Gu and Dao, 2023). In practice, SSMs are highly convenient. For example, in Vision Mamba (Zhu et al., 2024), the Mamba module replaces the Transformer module in the Vision Transformer (ViT) (Dosovitskiy et al., 2021), achieving better performance than ViT while also being more computationally efficient. Additionally, in the Video Mamba Suite (Chen et al., 2024), Mamba is evaluated as a plug-in or backbone, showing excellent results across 12 video understanding tasks, including action recognition and temporal action localization.
@@ 4.2. XRFMamba
Projection and Fusion. The Projection module is adapted from WiFiTAD (Liu et al., 2025) and comprises three sequential components: 1D Convolutional Operations, Group Normalization, and ReLU Activation. The Projection module standardizes the output channels of IMU and Wi-Fi data using convolutional operations, ensuring consistent channel dimensions for the subsequent fusion operation. The fusion of Wi-Fi and IMU data is performed through a weighted summation, as shown in the following equation:
@@
Here, $e_{wifi}$ and $e_{imu}$ represent Wi-Fi and IMU data outputs after projection, respectively. In our experiments, we observed that the IMU data performed better than Wi-Fi data. To maintain the stability of the fusion performance, we assigned a higher weight to the IMU data. Specifically, we set $\lambda=0.2$ in our experiments.
@@
Embedding. After fusion, the design of our embedding module takes inspiration from ActionMamba (Chen et al., 2024). In ActionMamba, the input to the Mamba backbone is not raw video but feature embeddings derived from the pre-trained model, InterVideo (Wang et al., 2022a). Following a similar approach, we introduce an embedding step prior to feeding the data into the Mamba backbone. However, due to the absence of widely accepted pre-trained models for IMU and Wi-Fi data in both academic and industrial contexts, a dedicated module is required to generate these embeddings. In this regard, we leverage the TSSE (Temporal Signal Semantic Embedding) module from WiFiTAD, which has proven effective in embedding Wi-Fi data for temporal action localization tasks. The TSSE embedding module is a dual-stream architecture, where one stream is based on the Transformer, which excels at learning global features, and the other is based on convolutional layers, which are effective at capturing local features. These two streams are fused at the output stage to create a comprehensive feature representation. The fused embeddings are then passed into the Mamba backbone for further processing.

### v2

@@ 3.1. Hardware Setups
Wi-Fi transceivers. We used ThinkPad X201 laptops equipped with Intel 5300 wireless network cards as the Wi-Fi transceivers. The Wi-Fi system consists of one transmitter and three receivers, positioned at the four corners of a 3.1 m×$\times$2.4 m rectangular area. This setup is inspired by deployment strategies used in several studies, such as Widar3.0 (Zheng et al., 2019), OneFi (Xiao et al., 2021), and XRF55 (Wang et al., 2024), which have shown that a multi-device configuration like this can enhance sensing performance. The devices are placed at a height of 120 cm to ensure effective coverage for full-body motion capture (Wang et al., 2019a). The transmitter operates with a single antenna, broadcasting packets at 200 per second on channel 128 (5.64 GHz). Each receiver is equipped with three antennas. Using the Wi-Fi CSI Tool (Halperin et al., 2011), we extract the channel state information (CSI) from 30 orthogonal frequency-division multiplexing (OFDM) subcarriers. The resulting CSI recordings are represented as a tensor of size $(200t)\times 1\times 3\times 3\times 30$, where $t$ is the recording duration in seconds. For example, for an 80-second continuous action, the CSI recordings yield a tensor of size $16000\times 1\times 3\times 3\times 30$, capturing detailed variations in the wireless channel over time for subsequent analysis.
@@
Azure Kinect: The Azure Kinect captures RGB, depth, and infrared images at 15 frames per second in 720P resolution using its toolkit. The Kinect is placed between the transmitter and the third receiver at a height of 115 cm to capture video footage of the movement of individuals in the scene. We can use open-source computer vision models to extract pose and mesh information from the videos, which will be useful for future tasks such as IMU- and Wi-Fi-based pose estimation and mesh reconstruction.
The experimental scenarios are shown in Fig. 2. Before data recording, all the aforementioned hardware devices undergo careful synchronization. Notably, XRF V2 selects commonly available consumer-grade products used in everyday household and personal settings, unlike XRF55 (Wang et al., 2024), which utilizes millimeter-wave radar and an RFID reader with a tag array.
@@
3.2. Action Sequence Proposal
We select three typical household scenarios for the experiment: the dining room, study room, and bedroom. These settings allow us to summarize users’ activities related to eating, studying, and sleeping, along with an associated series of actions. For example, activities such as taking medicine, drinking water, washing hands, and stretching serve as health-related indicators; and actions like opening curtains or turning on lights provide insights into interactions with household devices.
We communicate the above objective to the volunteers. Instead of assigning each person a fixed action sequence, we aim to increase the diversity of action sequences and incorporate personal habits relevant to the volunteers. Therefore, we ask each volunteer to recall their own daily action sequences in the dining room, study room, and bedroom, and estimate the execution time of each action in the sequence. Each volunteer recalls five action sequences based on the starting positions of the actions, with each sequence containing 7-10 actions. Fig. 4 shows examples of some action sequences, while Table 2 presents the proposed actions in each scene. In Appendix A, we demonstrate that these actions can be integrated with a large language model (LLM) assistant to enhance smart home intelligence.
@@ 3.3. Volunteer Coordination and Action Execution
We recruit 16 volunteers, with an age range of 22–34 years, a height range of 1.57–1.82 meters, and a weight range of 43–90 kgs. Volunteers can access the informed consent form online, which includes details about the study’s purpose, methods, participation duration, potential risks, and compensation.
Before performing the action sequences, we proportionally shorten the durations estimated by the volunteers to fall within 5–20 seconds. This adjustment is necessary because the originally estimated durations are often much longer; for instance, reading a book might be estimated as one hour. Performing such lengthy actions would significantly increase data collection time and complicate annotation. We also make reasonable adjustments to the execution durations several times, generating more action sequences. As a result, the action sequences in each scene become more diverse.
After that, we can adopt an automated annotation method to streamline the process. The action sequences proposed by the volunteers are converted into audio prompts using text-to-audio technology. These audio prompts are played in real time to guide the volunteers. Upon hearing an action instruction, the volunteer performs the corresponding action and transitions to the next action upon hearing the subsequent prompt. This process continues until the entire sequence is completed.
@@
In general, the duration of a single action sequence is approximately 50–80 seconds. After completing one action sequence, volunteers rest for 10 seconds, and after executing five sequences, they take a 5-minute break. These brief breaks provide the volunteers with proper rest, reduce fatigue, and maintain the overall quality of the data collection process. It also allows us to check the functionality of various devices and minimizes the cost of re-collecting data.
@@ 3.4. Dataset Statistics
Each of the 16 volunteers proposed 5 action sequences for each scene. By shortening durations and adjusting the length of individual actions, we generated multiple variants of these sequences. Specifically, in the bedroom, each volunteer performed three different variations of each action sequence, while in the study room and dining room, four different variations were performed for each sequence. Thus, XRF V2 dataset contains a total of: $16\times 5\times 3+16\times 5\times 4+16\times 5\times 4=880$ action sequences. After filtering out incomplete and erroneous sequences, we obtained 853 valid action sequences, with a total duration of 16 hours, 16 minutes and 8 seconds. These sequences were captured using synchronized Channel State Information (CSI) from four Wi-Fi transceivers, IMU data from two smartwatches, two smartphones, a pair of earbuds, and a pair of smart glasses, along with RGB+D+IR video streams from a Kinect camera. Furthermore, we implemented a stratified partitioning strategy maintaining an 80:20 ratio between training and testing sets across all environmental contexts and participants, resulting in 682 sequences (80%) allocated for model training and 171 sequences (20%) reserved for evaluation purposes. Table 3 shows detailed statistics of XRF V2 dataset.
@@
4. Background
State Space Models (SSMs), such as S4 (Gu et al., 2021), S5 (Smith et al., 2022), and Mamba (Gu and Dao, 2023), are a recent class of deep learning methods that have demonstrated advantages in both natural language processing (Gu and Dao, 2023) and computer vision (Zhu et al., 2024). Unlike Transformer-based models, which suffer from quadratic time complexity due to self-attention, SSMs operate with linear-time complexity, making them significantly more efficient for long sequence modeling. In action summarization, the input sensor sequences are much longer than those used in action recognition tasks, necessitating the ability to effectively model long-range dependencies. To this end, we base our approach on an advanced SSM model and propose the XRFMamba network.
@@ 4. Background
Figure 6. Basic Mamba Block (Gu and Dao, 2023).
Due to the flexibility of SSMs, they can switch to convolution during training and to recurrence during inference, providing fast performance in both stages. SSMs have a rich foundational knowledge, and introducing all aspects is beyond the scope of this paper. For more detailed information, please refer to Mamba (Gu and Dao, 2023) (structure shown in Fig. 6) . In practice, SSMs are highly convenient. For example, in Vision Mamba (Zhu et al., 2024), the Mamba module replaces the Transformer module in the Vision Transformer (ViT) (Dosovitskiy et al., 2021), achieving better performance than ViT while also being more computationally efficient. Additionally, in the Video Mamba Suite (Chen et al., 2024), Mamba is evaluated as a plug-in or backbone, showing excellent results across 12 video understanding tasks, including action recognition and temporal action localization.
5. Method
@@ 5.1. XRFMamba
(1) Projection and Fusion. The Projection module is adapted from WiFiTAD (Liu et al., 2025) and comprises three sequential components: 1D Convolutional Operations, Group Normalization, and ReLU Activation. The Projection module standardizes the output channels of IMU and Wi-Fi data using convolutional operations, ensuring consistent channel dimensions for the subsequent fusion operation. The fusion of Wi-Fi and IMU data is performed through a weighted summation, as shown in the following equation:
@@
Here, $e_{wifi}$ and $e_{imu}$ represent Wi-Fi and IMU data outputs after projection, respectively. In our experiments, we observed that the IMU data performed better than Wi-Fi data. To maintain the stability of the fusion performance, we assigned a higher weight to the IMU data. Specifically, we set $\lambda=0.2$ in our experiments. In this stage, the Projection module primarily ensures that the channel dimensions of the two data sources are consistent, preparing them for the subsequent fusion operation. The fusion module effectively combines the information from Wi-Fi and IMU, which is subsequently fed into further processing.
@@
(2) Embedding. After fusion, the design of our embedding module takes inspiration from ActionMamba (Chen et al., 2024). In ActionMamba, the input to the Mamba backbone is not raw video but feature embeddings derived from the pre-trained model, InterVideo (Wang et al., 2022a). Following a similar approach, we introduce an embedding step prior to feeding the data into the Mamba backbone. However, due to the absence of widely accepted pre-trained models for IMU and Wi-Fi data in both academic and industrial contexts, a dedicated module is required to generate these embeddings. In this regard, we leverage the TSSE (Temporal Signal Semantic Embedding) module from WiFiTAD, which has proven effective in embedding Wi-Fi data for temporal action localization tasks. The TSSE embedding module is a dual-stream architecture, where one stream is based on the Transformer, which excels at learning global features, and the other is based on convolutional layers, which are effective at capturing local features. These two streams are fused at the output stage to create a comprehensive feature representation. The fused embeddings are then passed into the Mamba backbone for further processing. The TSSE module learns the basic representations for both IMU and Wi-Fi data, overcoming the challenge of preparing representations for the Mamba backbone in the absence of pre-trained models for these modalities.